### v1

@@ 4.3.3 Anatomy Rendering (AR) Agent Functions
To ensure visual smoothness, $\mathcal{O}_{\text{AR}}$ interpolates the rotation and zoom transitions across frames. Zoom operations complete within three seconds, left/right/up/down rotation within seven seconds (three seconds movement, one second hold, 3 seconds return), and horizontal/vertical rotation within six seconds. The resulting overlays appear in the upper-right corner of the surgical video, allowing the surgeon to identify critical structures that require attention during the procedure.
4.4 Dataset
We construct a command dataset consisting of 240 user commands to evaluate our voice-directed SAOP. Among them, 44 commands belong to the IR agent, 81 commands correspond to the IV agent, and the remaining 115 commands are associated with the AR agent (Supplementary Table 3). This distribution reflects the diverse functional load of each agent. The AR agent accounts for nearly half of all commands because it offers a broader range of actions for 3D model interaction.
Each command also includes category annotations along three hierarchical dimensions: command structure, command type, and command expression (Supplementary Table 1). In the command structure dimension, single command contains one clearly defined action or data request, while composite command combines two or more actions or data requests. These composite command includes multi-action requests such as “move axial to the middle slice and zoom in”, and multi-data requests like “show physical and PFT info.”
The command type dimension classifies each command as explicit command, implicit command, or NLQ, reflecting the varying levels of linguistic clarity and ambiguity. Explicit command clearly state both the subject (the target data or agent) and the verb (the action), forming a grammatically complete and semantically clear instruction. These commands typically match the phrasing defined in the prompt instructions and do not require additional contextual reasoning such as “Show patient information”. Implicit command omit either the subject or the verb, making them grammatically incomplete and dependent on prior memory context. They often appear in continuous interactions where users skip information already mentioned in earlier commands. For example, users may omit the subject in “Zoom in”, “Initialize”, or “Rotate to the right”, or omit the verb as “Diagnosis” or “Surgical view”. The LLM should interpret such commands by leveraging previous memory states and contextual cues. NLQ take the form of questions, using conversational QA-style phrasing. Examples include “How old is the patient?”, “Can you magnify the sagittal view?”, and “Can you activate the right lung?”.
Finally, the command expression dimension characterizes lexical variations, indicating whether the command follows the baseline form, includes abbreviations, or paraphrases. Baseline command is the canonical form of expressions without any abbreviations or paraphrased expressions. Commands with abbreviations include medical terms such as PFT (Pulmonary Function Test) or RUL (Right Upper Lobe) that require domain knowledge, as well as general abbreviations like info (information). Commands with paraphrased expressions express the same intent as baseline commands but use alternative synonyms or phrasings. Because clinicians may prefer different terminology or phrasing styles, this category evaluates linguistic flexibility and contextual understanding. Examples include “Move front, front, front” (Coronal plus 30), “Coronal view to the middle slice” (Move coronal to 256), “Show from the front” (Show from the anterior), and “Rotate left and right” (Rotate horizontally).
@@ 4.4 Dataset
We additionally synthesize voice command datasets for four different speakers using Microsoft Edge TTS that we use in real-time speech function. Among the available voices, we select two female and two male speakers categorized under “News”, as they are native English speakers and provide clear pronunciation. Specifically, en-US-AriaNeural conveys polite and confident tones, en-US-JennyNeural sounds friendly, considerate, and comforting, en-US-GuyNeural expresses passion, and en-US-ChristopherNeural reflects reliability and authority.
@@
Patient data comes from the department of thoracic and cardiovascular surgery at Korea University Guro Hospital, involving a patient who had lung surgery using the Da Vinci robotic system. The patient data encompass surgical videos, clinical information, CT images, and 3D anatomical structure models. This study is ethically approved by the Institutional Review Board at Korea University Guro Hospital (IRB No. 2025GR0018), and informed consent was waived due to its retrospective design of the study.
5 Data availability
The command dataset (xlsx format) is provided in the Supplementary Data 1 file, and the synthesized voice command datasets (mp3 format) are available in the Github repository: https://github.com/helena-lena/SAOP. The de-identified patient data is not publicly available due to the privacy laws, specifically the Personal Information Protection Act of the Republic of Korea.
6 Code availability
The codes for the proposed voice-directed SAOP used in this study is available at the following GitHub repository: https://github.com/helena-lena/SAOP.
7 Acknowledgments
This research was supported by a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00436472). Additionally, this work was supported by the National Research Foundation of Korea (NRF) grant, funded by the Korea government (MSIT)(grant number: RS-2025-00518091).
8 Author contributions
H.P. implemented the system, conducted all experiments, analyzed the results, and drafted the manuscript. B.M.G., and J.H.L. collected the data and provided clinical feedback. B.H.C. managed the server for data storage and processing. S.K. advised on system implementation and evaluation, and contributed to manuscript preparation. H.K.K., and K.K. supervised the project from conceptualization to discussion, and contributed to manuscript preparation.
9 Competing interests
All authors declare no financial or non-financial competing interests.

### v3

@@ II-C3 AR Agent Functions
$\mathcal{O}_{\text{AR}}$ smoothly interpolates the rotation and zoom transitions, with zooms completing in three seconds, left/right/up/down rotations in seven seconds (three seconds movement, one second hold, three seconds return), and full horizontal/vertical rotations in six seconds. The resulting overlays appear in the upper-right corner of the surgical video, allowing the surgeon to identify critical anatomical structures during the procedures.
II-D Dataset
We construct a dataset of 240 user commands [^7] Data and code on https://github.com/helena-lena/SAOP: 44 for the IR agent, 81 for the IV agent, and 115 for the AR agent. This distribution reflects the differing functional loads of the agents, with the AR agent accounting for nearly half of all commands due to its broader range of actions and parameters. Each command also includes category annotations along three hierarchical dimensions: command structure, command type, and command expression, as summarized in Table I.
The command structure dimension categorizes each command as single or composite, indicating whether it requires decomposition into sequential sub-commands. The command type dimension classifies commands into explicit, implicit, and Natural Language Question (NLQ) forms, capturing different levels of linguistic clarity and ambiguity. Finally, the command expression dimension groups commands into baseline, abbreviation, and paraphrase to reflect lexical and phrasing variability, as clinicians express the same intent using different terminology or styles.
TABLE I: Hierarchical command category definitions and distributions.
Dimension Category Description and Examples Count
Command Structure Single Definition: Contains one clearly defined action or data request. 225
Example: “Show patient information.”, “Display CT images.”, “Remove models.”
Composite Definition: Combines two or more actions or data requests. 15
Example: “Move axial to the middle slice and zoom in.”, “Show physical and PFT info.”
Command Type Explicit Definition: States both subject (data/agent) and verb (action), and does not require contextual reasoning. 80
Example: “Show the anterior view.”, “Add the right upper lobe.”
Implicit Definition: Omits either the subject or the verb, making it dependent on the previous memory context. 80
Example: “Zoom in.”, “Initialize.”, “Surgical view.”, “Diagnosis.”, “Rotate to the right.”
NLQ Definition: Uses conversational or question-style phrasing. 80
Example: “How old is the patient?”, “Can you activate the right lung?”
Command Expression Baseline Definition: Expressions without any abbreviations or paraphrasing, matching phrasing defined in the prompt. 145
Example: “Display the pulmonary function test.”, “Coronal plus 256.”, “Show from the anterior.”
Abbreviation Definition: Includes abbreviated forms used in both medical terminology and general expressions. 15
Example: “Display the PFT info.”, “Remove RUL.”, “Show patient info.”
Paraphrase Definition: Expresses the same intent as baseline but uses alternative synonyms or phrasing styles. 80
Example: “Move front, front, front.”, “Move coronal to the middle slice.”, “Show from the front.”
@@ II-D Dataset
We additionally synthesize voice command datasets00footnotemark: 0 for four different speakers using Microsoft Edge TTS that we use in real-time speech function. Among the available voices, we select two female and two male speakers categorized under “News”, as they are native English speakers and provide clear pronunciation. Specifically, en-US-AriaNeural conveys polite and confident tones, en-US-JennyNeural sounds friendly, considerate, and comforting, en-US-GuyNeural expresses passion, and en-US-ChristopherNeural reflects reliability and authority.
@@
Patient data were obtained from the Department of Thoracic and Cardiovascular Surgery at Korea University Guro Hospital and include surgical videos, clinical information, CT images, and 3D anatomical models from a patient who underwent lung surgery using the da Vinci robotic system. This study was ethically approved by the Institutional Review Board at Korea University Guro Hospital (IRB No. 2025GR0018), and informed consent was waived due to its retrospective design.
III Evaluation Metrics
For each command, we evaluate the outcome of each workflow stage using a binary scheme (1: correct, 0: incorrect), as shown in Table II. STT checks whether the transcribed command, command correction (CC) evaluates the revised command, and command reasoning (CR) determines whether the LLM assigns the command to the appropriate agent. Agent function (AF) and agent parameters (AP) are outputs of AD function, which is correct only if both AF and AP are accurate. Orchestration flow (OF) assesses whether the workflow proceeds in the correct order, while invalid cycle (IC) is the number of times the workflow enters an invalid loop. Finally, we introduce the Multi-level Orchestration Evaluation Metric (MOEM), which comprehensively evaluates VISA from two perspectives: command-level orchestration performance and command-category performance.
The command-level orchestration evaluation measures two aspects: the stage-level accuracy that measures correctness at each workflow stage, and the workflow-level success rate (SR) that quantifies overall success throughout the entire workflow. The stage-level accuracy is the mean of binary outcomes across all commands for each stage:
where $N$ denotes the total number of commands, $i$ indicates the command index, and $o_{i,stage}$ is the binary result of the $i$-th command at a stage. The workflow-level SR evaluates each command under three conditions: strict, single-pass, and multi-pass as shown in Fig. 6. For each command $i$, a binary label $o_{i,condition}\in\{0,1\}$ indicates whether it satisfies the given success condition, and the $SR_{condition}$ is the average of these labels over all $N$ commands:
$\text{SR}_{\text{condition}}=\frac{1}{N}\sum_{i=1}^{N}o_{i,\text{condition}}.$ (17)
The command-category evaluation examines how structural and linguistic variations affect orchestration performance and identifies which categories are more prone to errors. For each category $c$, $SR_{multi-pass}$ is the mean of binary outcomes:
$\text{SR}_{multi-pass}(c)=\frac{1}{N_{c}}\sum_{i\in c}o_{i,\text{multi-pass}},$ (18)
where $N_{c}$ is the number of commands in category $c$. We also calculate the cross-category $SR_{multi-pass}$ for structure-type and type-expression pairs to explore how it jointly affects the performance. For each pair of categories $(c_{1},c_{2})$, the cross-category $SR_{multi-pass}$ is:
$\text{SR}_{\text{multi-pass}}(c_{1},c_{2})=\frac{1}{N_{c_{1},c_{2}}}\sum_{i\in(c_{1},c_{2})}o_{i,\text{multi-pass}},$ (19)
where $N_{c_{1},c_{2}}$ is the number of commands in the pair.
Figure 6: Success conditions for workflow-level SR, where green boxes represent success and red boxes denote success or fail: (a) In strict condition, the command is successful only if all stages succeed with no invalid loops; (b) In single-pass condition, it is also successful when intermediate failures are resolved in later stages, leading to a correct final outcome in a single attempt with no invalid loop; (c) In multi-pass condition, it is also successful when the final outcome is correct after multiple attempts, allowing up to three invalid loops.
TABLE II: Example of VISA workflow stage outcomes for each command.
Agent Command Structure Type Expression STT $C$ CR AF AP AD OF IC
IR Show patient information Single Explicit Baseline 1 1 1 1 1 1 1 0
IR Display the PFT info Single Explicit Abbreviation 1 1 1 1 1 1 1 0
IR Can you show pulmonary function test? Single NLQ Baseline 1 1 1 1 1 1 1 0
IR PFT results Single Implicit Abbreviation 1 1 1 1 1 1 1 0
IR lung function results Single Implicit Paraphrase 0 1 1 1 1 1 1 0
IR Show physical and PFT info Composite Explicit Abbreviation 1 1 1 1 1 1 1 0
IR Reset Single Implicit Baseline 1 1 1 1 1 1 1 0
IV Show the CT views Single Explicit Baseline 0 1 1 1 1 1 1 0
IV Coronal plus 100 Single Explicit Baseline 0 1 1 1 1 1 1 0
IV Sagittal minus 30 Single Explicit Baseline 1 1 1 1 1 1 1 0
IV Can you move CT forward? Single NLQ Baseline 1 1 1 1 1 1 1 0
IV Move front, front, front Single Implicit Paraphrase 0 1 1 1 1 1 1 0
IV Step to posterior Single Implicit Paraphrase 0 1 1 1 1 1 1 0
IV Axial zoom in Single Explicit Baseline 1 1 1 1 1 1 1 0
IV Can you get axial closer? Single NLQ Paraphrase 1 1 1 1 1 1 1 0
IV Minimize the axial image Single Explicit Paraphrase 1 1 1 1 1 1 1 0
IV Could you reduce coronal image? Single NLQ Paraphrase 0 1 1 0 0 0 1 0
IV Zoom out Single Implicit Baseline 1 1 1 1 1 1 1 0
IV Move axial to the middle slice and zoom in Composite Explicit Paraphrase 1 1 1 1 1 1 1 0
IV Can you move axial to 200 and coronal to 230? Composite NLQ Baseline 0 1 1 1 1 1 1 0
AR Show the 3D recon image Single Explicit Baseline 1 1 1 1 1 1 1 0
AR Can you load anatomical reconstruction? Single NLQ Paraphrase 1 1 1 1 1 1 1 1
AR Turn on RLL Single Explicit Abbreviation 1 1 1 1 1 1 1 0
AR Can you hide left lung? Single NLQ Paraphrase 1 1 1 1 1 1 1 0
AR Show anterior view Single Explicit Baseline 1 1 1 1 1 1 1 0
AR Can you look from the front? Single NLQ Paraphrase 1 1 1 1 1 1 1 0
AR Surgical view Single Implicit Baseline 1 1 1 1 1 1 1 0
AR Surgeon’s view Single Implicit Paraphrase 1 1 1 1 1 1 1 0
AR Rotate to the right Single Explicit Baseline 1 1 1 1 1 1 1 0
AR Would you rotate up? Single NLQ Baseline 1 1 1 1 1 1 1 0
AR Can you zoom in to RLL? Single NLQ Abbreviation 1 1 1 0 1 0 1 0
AR Zoom out Single Implicit Baseline 1 1 1 1 0 0 1 0
AR Can you zoom in and rotate to the left? Composite NLQ Baseline 1 1 1 1 1 1 1 0
AR Can you initialize and zoom in? Composite NLQ Baseline 1 1 1 1 1 1 1 0
AR Can you erase all? Single NLQ Paraphrase 1 1 1 1 1 1 1 0
All results available at https://github.com/helena-lena/SAOP.
IV Results
IV-A Command-level Orchestration Evaluation
IV-A1 Stage-level Accuracy
Figure 7: Stage-level accuracy results: (a) Overall stage-level accuracy; (b) Stage-level accuracy by task agent type; (c) Overall accuracy of action determination components; (d) Action determination component accuracy by task agent type; Error bars indicate 95% confidence intervals.
Figure 8: Workflow-level sucess rates results: (a) Overall results; (b) Results by task agent type; Error bars indicate 95% confidence intervals.
The VISA demonstrates strong error recovery capability from STT stage failures, maintaining high accuracy across subsequent stages as shown in Fig. 7 (a). The real-time audio stage accuracy is the performance of the wake-word detection model for the keyword “davinci”, which correctly recognized 266 out of 300 attempts ($88.6\%$). The remaining 34 attempts correspond to a false rejection rate of $11.3\%$, mainly due to the model misinterpreting “davinci” such as “Dabinci”, “David”, “Benchi”, or “dubbing”. For successful detections, the latency between speech detection and the “davinci” recognition averaged 1.97 seconds, showing smooth real-time interaction. Although STT accuracy is lower due to transcription errors, most are corrected in the CC stage. The CR stage maintains high accuracy in selecting the agent, AD accuracy shows a slight decline due to occasional mispredictions in the agent function or parameters. Finally, the OF achieves $100\%$ accuracy, confirming that all workflow stages execute in the correct order.
The stage-level accuracies across the three task-specific agents exhibit similar trends (Fig. 7b). The STT stage shows the largest variation, with the IV agent exhibiting the lowest accuracy due to frequent misrecognition of “CT” as “city”. Nevertheless, most recognition errors are corrected during the CC stage, the subsequent CR stage also maintains high performance across all agents, and the AD stage shows a slight decrease. Later stages generally show lower accuracy than the CC stage, reflecting the natural accumulation of prediction errors when performing stages step-by-step with an LLM. Interestingly, the IR agent even compensates for imperfectly corrected commands during the CR stage. For example, when the STT transcribes “age info” as “Age in full”, the CC stage refines it to “Age information in full”, allowing the CR stage to infer the intent as a request for the age column only. Overall, VISA demonstrates strong robustness, effectively mitigating early-stage errors and ensuring reliable agent execution.
In the AD stage, each task-specific agent determines the action uniquely mapped to an agent function, and its parameters specifying how the action should be executed. Our VISA achieves consistently high accuracy for the function and parameters, and because the mispredictions in the two components do not overlap, the combined AD accuracy is slightly lower (Fig. 7c). The accuracies across the three task agents exhibit similar trends, where the integrated AD accuracy is comparable to or slightly lower than the accuracies of function and parameter predictions (Fig. 7d). A slight decrease in accuracy from the IR to IV and AR agents is likely due to the greater variability and complexity of the commands in the latter tasks. Nevertheless, the differences remain marginal, indicating that the LLM effectively generalizes its reasoning capabilities across heterogeneous task agents.
IV-A2 Workflow-level Success Rate
The workflow-level SR improves progressively from strict to single-pass to multi-pass conditions as shown in Fig. 8 (a). The $SR_{strict}$ remains at 65.8%, primarily limited by transcription errors in the STT. In contrast, the $SR_{single-pass}$ increases to 89.2% as the LLM resolves early-stage failures through correction and reasoning. The $SR_{multi-pass}$ reaches 95.8%, showing that the workflow orchestrator agent returns to the real-time audio stage for invalid commands, allowing the user to restate the command more clearly. A similar trend appears in task-specific agents as shown in Fig. 8 (b). The IV agent shows the largest gap between strict and single-pass due to frequent STT errors, whereas the IR shows the smallest. As task complexity increases from IR to IV and AR, both $SR_{single-pass}$ and $SR_{multi-pass}$ decrease.
Figure 9: A sankey diagram illustrating recovery and failure paths of the command through stages from (a) STT, (b) $C$, (c) CR, and (d) AD.
IV-A3 Failure and Recovery Paths
We also analyze command processing paths from STT to AD stages as shown in Fig. 9. Most STT errors arise from confusing medical terms with phonetically similar non-medical words, such as recognizing “CT” as “city”, “coronal” as “corona”, and “lung” as “long”. Minor lexical errors also occur like recognizing “to” as “2”, “right” as “write”, “zoom” as “June”, and “add” as “at”. While most are resolved in the CC stage, a few remain, converting medical terms into non-medical words or nouns into adjective forms. CR failures occur when commands remain invalid after correction, whereas minor CC errors often do not affect agent selection, resulting in recovery. AD failures mainly arise from composite commands or mispredictions on functions or parameters. These findings highlight the need to support multi-step actions and refine prompts for challenging cases.
IV-B Command-category Evaluation
We evaluate $SR_{multi-pass}$ across hierarchical command categories and VISA robustly handles a wide range of structural and linguistic variations in user commands, as shown in Fig. 10 (a). Within the command structure category, composite commands are more challenging than single commands because the platform struggles to determine the correct agent function when multiple functions must be executed sequentially. For the command type category, the platform shows strong performance for explicit commands, implicit commands, and NLQ, demonstrating that the LLM can follow user intent with minimal performance degradation despite linguistic variations. Regarding the command expression category, commands using abbreviations achieve perfect performance because the LLM prompt explicitly includes the medical domain abbreviations. Commands with paraphrased expressions show slightly lower performance, as such variations are not included in the prompt and may introduce additional ambiguity.
Figure 10: Multi-pass success rate across command categories: (a) Multi-pass success rate by command category; (b) Multi-pass success rate by structure $\times$ type; (c) Multi-pass success rate by type $\times$ expression; Error bars in all plots indicate 95% confidence intervals.
Figure 11: Comparison by voice type: (a) Stage-level accuracy by voice type; (b) Success rate by voice type; Error bars are 95% confidence intervals.
We further analyze cross-category $SR_{multi-pass}$ as illustrated in Fig. 10 (b) and (c). In the structure–type relationship, single commands generally outperform composite ones, with composite failures occurring in three cases that require multiple actions, such as “Can you zoom in and rotate to the left?” or “Can you initialize and zoom in?”. All command types under the single structure maintain high performance, demonstrating that the VISA reliably processes straightforward, single-step commands. In the type–expression relationship, explicit commands achieve strong performance across all expression forms, whereas and only two failures arise from composite structure and entering invalid loops. In contrast, implicit commands are less robust to unseen synonyms or paraphrases, yet still maintain reasonable performance. Three failures are due to entering invalid loops, and one results from misinterpreting “pre-existing condition” and displaying both “comorbidities” and “diagnosis”. NLQs perform slightly better with paraphrase expressions because they usually contain richer contextual detail than short and implicit commands. Two failures appear: one when the VISA interprets “Could you reduce coronal image?” as adjusting slice position instead of zooming out, and another when it interprets “Can you view the superior angle?” as rotating upward rather than switching to the superior view. These results show that VISA generalizes well across different categories, with performance decreasing for composite or paraphrased commands that require multi-action and deeper reasoning on unseen expressions.
IV-C Applicability for Real Surgical Settings
In surgical environments, voice characteristics differ across speakers in tone, accent, speech rate, and pronunciation. To evaluate robustness to speaker variability, we compare four synthesized neural voices with human speech as shown in Fig. 11. Synthesized voices achieve higher STT accuracy due to their clear and native-like pronunciation, whereas human speakers recover more effectively from invalid loops by restating the command more clearly and emphasizing key information. Although some reasoning errors occur during the AD stage, all voice types maintain high stage-level accuracy. The workflow-level SR under strict, single-pass, and multi-pass conditions exhibit similar trends. Human speech exhibits lower $SR_{strict}$ because of STT failures, but ultimately achieves the highest $SR_{multi-pass}$ by recovering from invalid commands. These results demonstrate that VISA maintains stable performance across diverse voice conditions.
Furthermore, the operating room contains background noise and conversations from multiple surgeons and nurses. To ensure that wake-word “davinci” detection does not activate accidentally, we evaluate $15$ unsupported command examples as shown in Table III. The wake-word detection module remains active for one hour, and an unsupported command is spoken every four minutes. During one hour of testing, the VISA records a false alarm rate of 0, indicating no unintended activations. We also evaluate whether these $15$ unsupported commands are correctly classified as invalid after the wake word “davinci” is triggered. Three are misclassified as valid patient information requests, asking about vital sign, $CO_{2}$ pressure, and frozen section. Since these fields are not present in the given data columns, the VISA either overlays the previous video information or hides the overlay. Overall, our VISA reliably distinguishes task-related commands from unsupported commands, though a few cases still require refinement to prevent incorrect validation.
TABLE III: Results on unsupported commands
Example Wake-word Invalid
Replace the gauze. No Yes
Apply suction. No Yes
Clean the camera lens. No Yes
Wipe the third arm. No Yes
Switch between arm 1 and 2. No Yes
Prepare the hemostatic agent. No Yes
How long has the surgery been going on? No Yes
Is the patient’s vital sign stable? No No
What is the current CO2 pressure? No No
How long since the frozen section was sent? No No
The patient is coughing. No Yes
Prepare the stapler. No Yes
Prepare the ICG injection. No Yes
Prepare for the air leak test. No Yes
Insert the retrieval bag. No Yes
IV-D Analysis of Sequential Command Execution
We further evaluate whether VISA maintains continuity across sequential video clips [^8] Videos available at https://helena-lena.github.io/SAOP/ using its memory state, as shown in Fig. 12 and Fig. 13. The IV agent continuously updates axial, coronal, and sagittal slice positions within the predefined minimum and maximum constraints. For the AR agent, VISA preserves continuity in zoom operations by accurately computing the starting and ending scales while maintaining the minimum scale constraint of 1. The rotation and view adjustments also function correctly within the zoomed state, and the zoom out action follows the exact reverse path of zoom in. Additionally, when no command appears in the target video clip, the CC stage revises the empty input to “Select agent”, and the AD stage restores the previous agent state. For the “Reset” command, each agent restores the default settings of its function and the parameters defined in the prompt. These results demonstrate that VISA preserves temporal and functional consistency across video clips, ensuring stable and continuous state transitions.
V Discussion
We demonstrate the effectiveness of the hierarchical multi-agent framework with LLM-based orchestration. The workflow orchestrator agent reliably plans and selects appropriate functions based on the current state and predefined decision rules. The hybrid design of LLM-driven reasoning with probability-based decision rules resolves invalid loops and enables workflow stages to progress smoothly. Although computing probabilities for all functions slows inference compared to direct function prediction, it could enable conditional probability–based planning to handle increasingly complex paths as the number of agents and functions grows. Separating workflow functions from task-specific agent functions further improves modularity and scalability, allowing flexible integration of new agents or functions. Expanding clinically useful agents, such as detecting critical blood vessels, supporting intraoperative navigation, and assisting in robotic planning, will enhance the clinical applicability.
While our VISA demonstrates the feasibility of LLM-based orchestration for surgical assistance, the observed errors reveal opportunities for improvement. At the STT stage, most errors occur with medical domain-specific terminology, suggesting the need to fine-tune lightweight STT models on medical terms to improve transcription accuracy. In the CC stage, the LLM revises misrecognized words to their intended medical terms, yet augmenting text correction rules or fine-tuning a correction model on medical context could further enhance robustness in failure cases. At the AD stage, most errors arise from composite commands, indicating the need to extend the VISA to support multi-step sequential operations. Additionally, since the current prompting strategy are closely coupled with the Gemma3 [22] model, performance may vary across different LLMs. Developing a self-evolving orchestration mechanism that learns from prior errors and dynamically refines its prompts could ensure adaptability and consistent performance across models in future implementations.
From a dataset perspective, the evaluation uses 240 commands, including both human and synthesized voice inputs from four speakers to simulate variability. However, the dataset does not fully represent real clinical variability in phrasing, terminology, or multilingual mixing observed in real clinical settings. Future work may involve fine-tuning models on multilingual medical data or using fast and commercially available multilingual STT systems. Moreover, the clinical information displayed by the IR agent is based on a predefined set of columns selected by clinicians. However, the required fields may vary across departments and individual surgeons, indicating the need for customizable data configurations. For 3D anatomical models displayed by AR agent, using commercial or clinically validated software could provide more precise anatomical structures and support more sophisticated manipulation for clinical use.
Regarding computational efficiency, we employ Gemma3 27B (gemma3:27b-it-qat) as the core LLM for all VISA stages, a quantization-aware trained version that preserves high performance with reduced memory usage. We set the temperature to 0 to minimize randomness, use num_predict of 1-1 for unrestricted generation, and fix the seed at 42 for reproducibility. The model runs locally via Ollama [^9] https://github.com/ollama/ollama and occupies approximately 17 GB when preloaded onto a single NVIDIA RTX 4500 Ada GPU (24 GB), while the second GPU handles other APIs, models, or agents. However, cumulative latency from sequential LLM calls remains a challenge in time-sensitive surgical environments. This delay becomes more pronounced when the LLM engages in complex reasoning or generates longer output. Latency may be reduced by limiting LLM reasoning to essential stages, caching reasoning patterns, shortening outputs, and adapting to surgeon-specific interaction patterns. Lastly, for 3D model rendering, increasing mesh decimation or frame skipping improves speed but may reduce smoothness during rotation and zoom.
Figure 12: Video examples of consecutive scenarios for the IV agent.
Figure 13: Video examples of consecutive scenarios for the AR agent.